OpenStudy (anonymous):

a perimeter of a rectangle is 42 ft the length is 7ft longer than the width. Find the dimensions write a system of linear equaton and solve the system Let x be length an y be width 2x+2y= what

OpenStudy (anonymous):

? ur question is nt clear

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

Let x be length an y be width 2x+2y= what?

OpenStudy (anonymous):

I do not know

OpenStudy (anonymous):

wouldit be 4

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

sorry i was out..il help u

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

 so u can set up an equation x+x+x+7+x+7 = 42 4x + 14 = 42 x = 7